Colorado Corporation produces and sells waste baskets for home and office use. Its manufacturing plant has a capacity to produce 12,000 baskets each month. Current production and sales are 10,000 baskets. The company normally charges $30 per basket and, as shown below, reports an operating income of $50,000 each month. Unit cost information and monthly income statement are presented below: Per unit Total for 10,000 units Sales revenues $30.00 $300,000 Variable expenses: Manufacturing (DM, DL & VFOH) $13.00 $130,000 Selling expenses 4.00 40,000 Fixed expenses: Fixed manufacturing overhead 1.00 10,000 Fixed selling and administrative 7.00 70,000 Total costs $25.00 250,000 Operating income $ 50,000 8. Minnesota, Inc., has contacted Colorado about purchasing 2,000 baskets at a discounted price. For this one-time special order, what is the minimum acceptable price that Colorado should accept, assuming no variable selling expenses will be incurred? A) $25.00 B) $17.00 C) $13.00 D) $30.00 9. Minnesota, Inc., has contacted Colorado about purchasing 2,000 baskets at $20.00 each. If this one-time special order is accepted, current sales and total fixed expenses would not be affected. In addition the company will not incur variable selling expenses (sales commissions) on this order. If Colorado accepts the special order, its operating profits will be A) $22,000. B) $20,000. C) $50,000. D) $64,000.
Answer #1

8) Minimum acceptable price is the variable cost per unit. so in this question Variable manufacturing cost will be minimum acceptable price because there is no variable selling expense

so answer is c) $13.00

9) Operating profit = Sales-Variable manufacturing cost

= (2000*20)-(2000*13)

Operating profit from special order = $14000

Colorado's operating profit = 14000+50000 = $64000

so answer is d) $64000